Write the equation of the line, given:
a. m= -1/2, though (2,4)
b. Through (1,5) & (-2,1)

Answer by uma(370)

(1) Given m= - 1/2
The line passes through (2,4)
Equation of the line with slope -1/2 and through (2,4) is given by
y - 4 = -1/2 (x - 2)
==> 2y - 8 = -(x - 2)
==> 2y - 8 = -x + 2
==> x + 2y = 10 is the required equation.
(2) Line through (1,5) and (-2,1)
Equation is (y-5)/(x-1) = (1-5)/(-2-1)
==> (y-5)/(x-1) = -4/-3
==> -3(y-5) = -4(x-1)
==> -3y + 15 = -4x + 4
==> 4x -3y + 11 = 0 is the required equation.